OpenStudy (maeganr):

Question: One drink contains 5% fruit juice and another contains 50% fruit juice. How many gallons of each type should be combined to make 5 gallons of a drink that contains 40% fruit juice? PLEASE HELP ASAP. I WILL GIVE OUT MEDALS AND FAN YOU. (:

OpenStudy (maeganr):

@zaynahf @CalebBeavers @JuanitaM

OpenStudy (anonymous):

i so drew a blank...:)

OpenStudy (maeganr):

I dont get it. Im so confused.

OpenStudy (anonymous):

i think i got it.

OpenStudy (maeganr):

Can you help me?

OpenStudy (anonymous):

x of .50 juice 5-x of .05 juice amount of juice in final is .50x + .05(5-x) this equals .40 of 10 = 4 so, .50x + .05(5-x) = 4 now solve for x

OpenStudy (anonymous):

oop ,, meant to be 5 not 10

OpenStudy (maeganr):

What?........

OpenStudy (anonymous):

.40 of 5 = 2 so .50x + .05(5-x) = 2

OpenStudy (anonymous):

let me try to explain... you have a total of 5 gallons of juice part 50% juice (.50x) the other part 5% juice .05(5-x)

OpenStudy (maeganr):

ok

OpenStudy (anonymous):

now that solution must equal 40% juice...so that is why you have to multiple .40*5 =2

OpenStudy (anonymous):

now set both sides to equal each other.... .50x +.25 - .05x = 2 .45x + .25 = 2 .45x = 2 - .25 .45x = 1.75 x = 3.9 (rounded)....do you get this?

OpenStudy (maeganr):

Sorta....

OpenStudy (anonymous):

check this website out....maybe reading their explanations may help you get it. http://www.purplemath.com/modules/mixture2.htm

OpenStudy (anonymous):

to finish, i would use x = 4 which tellls you that you will use 4 gal of 50% juice thus leaving 6 gallons of 5% juice Good luck...

OpenStudy (maeganr):

Im gonna just write that and hope for the best.. lol thanks(:

OpenStudy (anonymous):

LOL.... you will do fine, mixtures can get a little confusing...